OpenStudy (anonymous):

What is the volume, in cubic inches, of Type 1 containers?

OpenStudy (anonymous):

Containers Radius (in inches) Volume (in cubic inches) Type 1 3 ? Type 2 4 113.11

OpenStudy (anonymous):

@DemolisionWolf can u help me :) haha

OpenStudy (anonymous):

haha sure, I really should log off and do my own homework soon!!!

OpenStudy (anonymous):

so type 1 has a radius of 3 and a volume of ? and type 2 has a radius of 4 and a volume of 113.11? am I right so far?

OpenStudy (anonymous):

so what we have to figure out is what type of shape is the container of type 2. is it a cylinder, cone, sphere. so we will use the formulas for those shapes and see if we can figure out what shape it is. if the radius for type 2 is 4 and the volume is 113.11 then lest check: cylinder check: volume = pi*r^2 * h pi*4^2 *h = 113.11 [T/F?] sphere check: volume = (4/3)pi*r^3 113.11 = (4/3)pi*4^3 [T/F?] cone check: volume = (1/3)pi*r^2 *h 113.11 = (1/3)pi*4^2 *h [T/F?]

OpenStudy (anonymous):

Hmm Im lost